# Neutrinoless $\beta\beta$ decay nuclear matrix elements complete up to N2LO in heavy nuclei

### v1

@@ Abstract
We evaluate all nuclear matrix elements (NMEs) up to next-to-next-to leading order (N2LO) in chiral effective field theory ($\chi$EFT) for the neutrinoless double-beta ($0\nu\beta\beta$) decay of the nuclei most relevant for experiments, including 76Ge, 100Mo, and 136Xe. We use the proton-neutron quasiparticle random-phase approximation (pnQRPA) and the nuclear shell model to calculate the N2LO NMEs from very low-momentum (ultrasoft) neutrinos and from loop diagrams usually neglected in $0\nu\beta\beta$ studies. Our results indicate that, separately, the new N2LO NMEs can have up to a $15\%$ impact on the total NME, but cancellations lead to an overall N2LO contribution centered around $-5\%$ for the shell model and $+10\%$ for the pnQRPA, with sizable uncertainties due to the short-range nature of the loop NMEs. The sign discrepancy between many-body methods is common to all studied nuclei and points to the different behaviour of the intermediate states of the $0\nu\beta\beta$ decay. In addition, our results support the $\chi$EFT prediction that the ultrasoft NME is closely related to contributions usually referred to as “beyond the closure approximation”.
@@ 1 Introduction
Exploring fundamental physics from $0\nu\beta\beta$ experiments requires the knowledge of the nuclear matrix elements (NMEs) governing the decay. However, they currently form a major obstacle as even the best NME calculations [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] carry large uncertainties [1, 34]. Moreover, a recent $0\nu\beta\beta$ derivation using chiral effective field theory ($\chi$EFT) [35] has pointed out additional uncertainties, showing that the so-called short-range NME, originally proposed at next-to-next-to-leading order (N2LO) in the power counting, had to be promoted to leading order (LO) for the theory to be renormalizable [36, 37]. Indeed, the LO short-range NME is in the ballpark of a 50% correction for various many-body methods: the nuclear shell model (NSM) [38], the proton-neutron quasiparticle random-phase approximation method (pnQRPA) [38], the ab initio valence-space in-medium similarity renormalization group (VS-IMSRG) approach [19, 18], and the generalized contact formalism that combines the NSM with quantum Monte Carlo (QMC) methods [20].
@@ 2 Neutrinoless Double-Beta Decay
×$\times\left|M_{\rm L}^{0\nu}+M_{\rm S}^{0\nu}+M_{\rm usoft}^{0\nu}+M_{\rm loops}^{0\nu}\right|^{2}\;,$ (1)
where $G_{0\nu}$ is a phase-space factor for the electrons [43], $g_{\rm A}\simeq 1.27$ is the axial-vector coupling to the nucleon, and $m_{\beta\beta}=\sum_{i}U_{ei}^{2}m_{i}$ (normalized to the electron mass $m_{e}$), which characterizes the lepton-number violation, depends on the neutrino masses $m_{i}$ and mixing matrix $U$. Here we focus on the NMEs: the LO long- and short-range, $M_{\rm L}^{0\nu}$ and $M_{\rm S}^{0\nu}$ [38], and the N2LO terms: $M_{\rm usoft}^{0\nu}$ coming from ultrasoft neutrinos and $M_{\rm loops}^{0\nu}$ resulting from loop diagrams that cannot be absorbed into the nucleon form factors entering $M^{0\nu}_{\rm L}$.
@@
$M^{0\nu}_{\rm loops}$ $=M^{0\nu}_{\rm loops,VV}+M^{0\nu}_{\rm loops,AA}$
$+M^{0\nu}_{\rm loops,US}+M^{0\nu}_{\rm loops,CT}\;,$ (7)
with each part ($X$=VV, AA, US, CT) given by
$M^{0\nu}_{{\rm loops},X}=\frac{4R}{\pi g_{\rm A}^{2}}\langle 0^{+}_{f}||\sum_{m,n}\tau^{-}_{m}\tau^{-}_{n}\mathcal{V}^{(m,n)}_{X}||0^{+}_{i}\rangle\;.$ (8)
@@
$-f_{0}(\hat{q})-24g^{2}_{A}F^{2}_{\pi}C_{T}(L_{\pi}+1)\Bigg{)}j_{0}(qr)\;,$ (10)
and the counterterm (CT)—that should absorb the renormalization scale ($\mu$) dependence of the VV, AA, and US terms—is given by
$\begin{split}&\mathcal{V}^{(m,n)}_{\rm CT}=\int{\rm d}q\;q^{2}e^{-q^{2}/(2\Lambda^{2})}\frac{g^{2}_{A}\hat{q}}{3(4\pi F_{\pi})^{2}}\\ &\times\Big{(}-\frac{5}{6}g^{\pi\pi}_{\nu}\frac{\hat{q}}{(1+\hat{q})^{2}}+g^{\pi N}_{\nu}\frac{1}{1+\hat{q}}\Big{)}\\ &\times(j_{0}(qr)\bm{\sigma}_{m}\cdot\bm{\sigma}_{n})-j_{2}(qr)S_{mn}^{\rm T})\;.\end{split}$ (11)
In this work we follow [35] and use the ultrasoft scale $\mu_{\text{us}}=m_{\pi}$. Since the ultrasoft (US) loop term $M^{0\nu}_{\rm loops,US}$ is proportional to $\ln\frac{\mu_{\rm us}}{m_{\pi}}$, it vanishes with this choice. We use the short-hand notations $\hat{q}=q^{2}/m_{\pi}^{2}$, $L_{\pi}=\ln(\mu^{2}/m_{\pi}^{2})$, the $f_{n}(\hat{q})$ functions from [35], $C_{T}=\pi/m_{\rm N}(1/(a_{t}^{-1}-\mu)-1/(a_{s}^{-1}-\mu))$ with $a_{s}^{-1}=-8.3$ MeV and $a_{t}^{-1}=36$ MeV, $g_{\nu}^{\pi\pi}=-10.84-36/5\ln(\mu/m_{\rho})$ from the average of the lattice QCD evaluations [49, 50], $F_{\pi}=92.28$ MeV and $g_{\nu}^{\pi N}$=0. Since the constants $C_{T}$ and $g_{\nu}^{\pi\pi}$ are evaluated around the $\rho$ mass, we vary the scale $\mu$ in the range $\mu=500\cdots 1500$ MeV as in [19]. We regularize the loop NMEs with the same Gaussian regulators as the LO short-range NME.
Table 1: Central values and ranges of all NMEs up to N2LO: LO long range (second and sixth columns), LO short range (third and seventh), N2LO ultrasoft (fourth and eighth) and N2LO loops (fifth and ninth). The NME ranges cover two interactions for the NSM or a range of $g_{pp}$ values for the pnQRPA, as well as various couplings $g_{\nu}^{\rm NN}$, Gaussian regulators $\Lambda=349\cdots 550$ MeV, CD-Bonn and Argonne SRCs, and the renormalization scale range $\mu=500\cdots 1500$ MeV (see text for details).
@@
L S usoft loops L S usoft loops
@@ 4.1 Ultrasoft NMEs and closure approximation
Table 1 presents the results of all NMEs up to N2LO calculated with the NSM and pnQRPA. The central values and ranges cover the interactions, couplings, regulators and SRCs discussed in Secs. 2 and  3. Focusing on the central values, the results in Table 1 indicate that for the NSM the ultrasoft NME is about $5\%$ and has opposite sign compared to the total LO NME—summing the long- and short-range parts. The only exception is 48Ca, where the ultrasoft NME is suppressed. In contrast, for the pnQRPA the ultrasoft NME has the same sign and $5\%-10\%$ of the value of the LO NME. Nonetheless, in 100Mo it reaches $20\%$ and it is very small in 116Cd and 136Xe.
In order to investigate the opposite sign of the ultrasoft NME in NSM and pnQRPA calculations, Fig. 1 compares the running sums of $M^{0\nu}_{\rm usoft}$ in 76Ge as a function of the excitation energy of the intermediate states, $E$, for both many-body methods. The results cover the NSM interactions and pnQRPA $g_{\rm pp}$ range presented in Sec. 3, complemented with two additional NSM interactions. Figure 1 shows that up to $E\approx 8$ MeV the two running sums are rather similar. However, they differ markedly around $8\text{ MeV}\lesssim E\lesssim 13\text{ MeV}$, where large cancellations and a resulting sign change appear in the pnQRPA but are not present in the NSM results. In fact, this behaviour is common to all the studied nuclei with the pnQRPA, while in the NSM we only observe $M^{0\nu}_{\rm usoft}$ cancellations, although milder, in 48Ca. Similar differences between running sums in the NSM and pnQRPA have been noticed in the context of $2\nu\beta\beta$ decay [63]—which involves a similar sum over $1^{+}$ intermediate states. While the NSM may not capture these high-energy contributions because of missing spin-orbit partners in the valence spaces used, the pnQRPA may overestimate them. Measurements of charge-exchange reactions up to the GT resonance energy in both $\beta^{-}$ and $\beta^{+}$ directions, currently limited to lower energies [64, 65], could solve this discrepancy.
@@
Figure 1: Running sum of the ultrasoft NME of 76Ge, $M^{0\nu}_{\rm usoft}$, as a function of the excitation energy of the intermediate $1^{+}$ states, $E$, for the NSM (red) and pnQRPA (blue). The NSM results include the JUN45 (solid), GCN2850 (dashdotted), RG.prolate [66] (dotted) and JJ4B [67] (dashed), while the pnQRPA results comprise $g_{\rm pp}$ values adjusted to $2\nu\beta\beta$-decay with $g_{\rm A}^{\rm eff}=1.0$ (dotted) and $g_{\rm A}=1.27$ (dashed).
@@
In $\chi$EFT [35], $M^{0\nu}_{\rm usoft}$ can be seen as a correction beyond closure [68, 69, 70, 71, 72, 73, 74], since it is the lowest-order NME that explicitly depends on the structure of the intermediate states. To test this prediction, we compare the difference of $M^{0\nu}_{\rm L}$ calculated with and without closure approximation, $\Delta_{\rm cl}\equiv M^{0\nu}_{\rm non-cl}-M^{0\nu}_{\rm L}$, with $M^{0\nu}_{\rm usoft}$. We perform all calculations for the $\Delta_{\rm cl}$ in the pnQRPA and use the results of [72, 73, 74] for the NSM—for 130Te we average between the closure energies for 136Xe [73] and 124Sn [74]. Figure 2 shows the comparison. For the NSM, $M^{0\nu}_{\rm usoft}$ is actually a very good approximation for beyond closure effects in all nuclei, strongly supporting the $\chi$EFT prediction. In contrast, for the pnQRPA, $M^{0\nu}_{\rm usoft}$ amounts to roughly twice $\Delta_{\rm cl}$ in all nuclei, except for 136Xe where the two quantities are small. Nonetheless, the sign of both quantities is consistently the same. $M^{0\nu}_{\rm usoft}$ may overestimate beyond closure effects in the pnQRPA results because the latter receive significant contributions from intermediate states with different total angular momentum than $J=1$ [45, 44]. Another difference is that the cancellations in $M^{0\nu}_{\rm usoft}$ highlighted in Fig. 1 are much milder for $M^{0\nu}_{\rm non-cl}$ [59].
@@
Figure 2: Difference between non-closure and LO long-range NMEs ($\Delta_{\text{cl}}$, orange and light blue bars) compared with the ultrasoft NME ($M^{0\nu}_{\text{usoft}}$, read and dark blue bars) obtained by the NSM and QRPA for all decays studied in this work. The NSM results use closure energies from [72, 73, 74] (solid bars) except for 130Te where we average the closure energies of 124Sn and 136Xe (open bar). Results for Argonne SRCs.
@@ 4.2 N2LO loop NMEs
Table 1 presents the calculated central values and ranges of the N2LO loop NMEs. While the central values are typically small, less than $5\%$ of the total LO NME for both many-body methods, the uncertainties are significant. When included, $M^{0\nu}_{\rm loops}$ can reach up to $10\%$ of the total LO NME for the pnQRPA—$15\%$ for 100Mo—, and up to $5\%$ for the NSM, which are roughly the relative effects of the ultrasoft NMEs.
Figure 3: Radial distributions of the N2LO loop NME for 76Ge, $C^{0\nu}_{\rm loops}$, obtained with the pnQRPA (left) and NSM (right). The results cover Gaussian regulators $\Lambda=349$ MeV (red) and $\Lambda=550$ MeV (blue), renormalization scales $\mu=500$ MeV (dark colors) and $\mu=1500$ MeV (light), and Argonne (dashed lines) and CD-Bonn (solid) SRCs.
To illuminate the origin of the large uncertainties for $M^{0\nu}_{\rm loops}$, Fig. 3 shows, for the NSM and pnQRPA, their radial distribution for 76Ge, which satisfies $M^{0\nu}_{\rm loops}=\int_{0}^{\infty}C^{0\nu}_{\rm loops}(r){\rm d}r$. Different curves show the results for different regulators $\Lambda$, scales $\mu$, and SRCs. Qualitatively, the radial distributions look very similar in the pnQRPA (left) and NSM (right), but the overall scales are quite different. This is consistent with the comparison of the LO NMEs [38]. Figure 3 also indicates that the main sensitivity of the radial distribution of $M^{0\nu}_{\rm loops}$ is to the regulator $\Lambda$ (difference between the blue and red curves: choosing a dipole regulator like in [42] gives results in between these two results). This sensitivity stems from the short-range—high-momentum—nature of the loop NME, also highlighted in Fig. 3, which concentrates at short-distances almost like the LO short-range NME [38]. This makes the N2LO loop NME very sensitive to SRCs as well: choosing either CD-Bonn or Argonne parameterizations can even change the sign of $M^{0\nu}_{\rm loops}$. In contrast, the dependence on the scale $\mu$ is rather mild.
@@
124Sn 4.88 2.18 0.58 -0.49 0.19 0.85 3.20 0.90 -0.26 -0.11 0.05 0.24
130Te 3.36 1.52 0.44 -0.39 0.15 0.67 3.24 0.94 -0.27 -0.12 0.06 0.19
136Xe 2.71 0.99 0.11 -0.23 0.09 0.40 2.60 0.75 -0.22 -0.09 0.04 0.15
Table 2 gives a decomposition of the VV, AA and CT components of the N2LO loops NME for a particular combination of nuclear interaction, couplings, regulators and SRCs—common for the NSM and pnQRPA results. The LO and N2LO ultrasoft NMEs obtained with the same combination of parameters are also listed for comparison. The results for the components of $M^{0\nu}_{\rm loops}$ indicate that, both for the NSM and pnQRPA, the different parts cancel, with the AA and CT terms sharing the same sign as the LO NMEs, and VV carrying systematically the opposite sign. Individually, each of the three N2LO loops contributions can be of similar size, or even larger than the N2LO ultrasoft NME; however, due to the cancellation, the total $M^{0\nu}_{\rm loops}$ tends to be smaller than $M^{0\nu}_{\rm usoft}$, see Table 1.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Exploring fundamental physics from $0\nu\beta\beta$ experiments requires the knowledge of the nuclear matrix elements (NMEs) governing the decay. However, they currently form a major obstacle as even the best NME calculations [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] carry large uncertainties [1, 35]. Moreover, a recent $0\nu\beta\beta$ derivation using chiral effective field theory ($\chi$EFT) [36] has pointed out additional uncertainties, showing that the so-called short-range NME, originally proposed at next-to-next-to-leading order (N2LO) in the power counting, had to be promoted to leading order (LO) for the theory to be renormalizable [37, 38]. Indeed, the LO short-range NME is in the ballpark of a 50% correction for various many-body methods: the nuclear shell model (NSM) [39], the proton-neutron quasiparticle random-phase approximation method (pnQRPA) [39], the ab initio valence-space in-medium similarity renormalization group (VS-IMSRG) approach [19, 18], and the generalized contact formalism that combines the NSM with quantum Monte Carlo (QMC) methods [20].
@@ 2 Neutrinoless Double-Beta Decay
×$\times\left|M_{\rm L}^{0\nu}+M_{\rm S}^{0\nu}+M_{\rm usoft}^{0\nu}+M_{\rm loop}^{0\nu}\right|^{2}\;,$ (1)
where $G_{0\nu}$ is a phase-space factor for the electrons [44], $g_{\rm A}\simeq 1.27$ is the axial-vector coupling to the nucleon, and $m_{\beta\beta}=\sum_{i}U_{ei}^{2}m_{i}$ (normalized to the electron mass $m_{e}$), which characterizes the lepton-number violation, depends on the neutrino masses $m_{i}$ and mixing matrix $U$. Here we focus on the NMEs: the LO long- and short-range, $M_{\rm L}^{0\nu}$ and $M_{\rm S}^{0\nu}$ [39], and the N2LO terms: $M_{\rm usoft}^{0\nu}$ coming from ultrasoft neutrinos and $M_{\rm loop}^{0\nu}$ resulting from loop diagrams that cannot be absorbed into the nucleon form factors entering $M^{0\nu}_{\rm L}$.
@@
$M^{0\nu}_{\rm loop}$ $=M^{0\nu}_{\rm loop,VV}+M^{0\nu}_{\rm loop,AA}+M^{0\nu}_{\rm loop,CT}$
$+M^{0\nu}_{\rm loop,usoft}\;,$ (7)
with each part ($X$=VV, AA, CT, usoft) given by
$M^{0\nu}_{{\rm loop},X}=\frac{4R}{\pi g_{\rm A}^{2}}\langle 0^{+}_{f}||\sum_{m,n}\tau^{-}_{m}\tau^{-}_{n}\mathcal{V}^{(m,n)}_{X}||0^{+}_{i}\rangle\;.$ (8)
@@
$-f_{0}(\hat{q})-24g^{2}_{A}F^{2}_{\pi}C_{T}(L_{\pi}+1)\Bigg{)}j_{0}(qr)\bigg{]}\;,$ (10)
the counterterm (CT)—that should absorb the renormalization scale ($\mu$) dependence of the VV, AA, and usoft terms—is given by
$\begin{split}&\mathcal{V}^{(m,n)}_{\rm CT}=\int{\rm d}q\;q^{2}e^{-q^{2}/(2\Lambda^{2})}\frac{g^{2}_{A}\hat{q}}{3(4\pi F_{\pi})^{2}}\\ &\times\Big{(}-\frac{5}{6}g^{\pi\pi}_{\nu}\frac{\hat{q}}{(1+\hat{q})^{2}}+g^{\pi N}_{\nu}\frac{1}{1+\hat{q}}\Big{)}\\ &\times(j_{0}(qr)\bm{\sigma}_{m}\cdot\bm{\sigma}_{n}-j_{2}(qr)S_{mn}^{\rm T})\;,\end{split}$ (11)
and the ultrasoft term—which, if one uses the leading-order $\chi$EFT Hamiltonian, cancels the dependence on the ultrasoft scale ($\mu_{\rm us}$) of $M_{\rm usoft}^{0\nu}$—is defined as
$\begin{split}&\mathcal{V}^{(m,n)}_{\rm usoft}=\int{\rm d}q\;q^{2}e^{-q^{2}/(2\Lambda^{2})}\bigg{[}-\frac{2g^{4}_{A}}{3(4\pi F_{\pi})^{2}}\frac{{q}^{2}}{{q}^{2}+m_{\pi}^{2}}\\ &\times(j_{0}(qr)\bm{\sigma}_{m}\cdot\bm{\sigma}_{n}-j_{2}(qr)S_{mn}^{\rm T})\\ &-\frac{2g^{4}_{A}}{(4\pi F_{\pi})^{2}}\frac{{q^{2}}}{{q^{2}}+m^{2}_{\pi}}+\frac{g^{2}_{A}}{(4\pi)^{2}}48C_{T}\bigg{]}\ln\left(\frac{m^{2}_{\pi}}{\mu_{\text{us}}^{2}}\right)\;.\\ \end{split}$ (12)
We vary the ultrasoft scale range as $\mu_{\text{us}}=m_{\pi}/2\dots 2m_{\pi}$, given that $M^{0\nu}_{\rm loop,usoft}$ is proportional to $\text{ln}(m^{2}_{\pi}/\mu_{\text{us}}^{2})$. We use the short-hand notations $\hat{q}=q^{2}/m_{\pi}^{2}$, $L_{\pi}=\ln(\mu^{2}/m_{\pi}^{2})$, the $f_{n}(\hat{q})$ functions from [36], $C_{T}=\pi/m_{\rm N}(1/(a_{t}^{-1}-\mu)-1/(a_{s}^{-1}-\mu))$ with $a_{s}^{-1}=-8.3$ MeV and $a_{t}^{-1}=36$ MeV, $g_{\nu}^{\pi\pi}=-10.84-36/5\ln(\mu/m_{\rho})$ from the average of the lattice QCD evaluations [50, 51], $F_{\pi}=92.28$ MeV and $g_{\nu}^{\pi N}$=0. Since the constants $C_{T}$ and $g_{\nu}^{\pi\pi}$ are evaluated around the $\rho$ mass, we vary the scale $\mu$ in the range $\mu=500\cdots 1500$ MeV as in [19]. We regularize the loop NMEs with the same Gaussian regulators as the LO short-range NME.
In this work, we give results combining the two N2LO terms that depend on $\mu_{\rm us}$, which we refer to as total ultrasoft NME, $M_{\rm tot,usoft}^{0\nu}$. On the other hand, we label the sum of the remaining loop NMEs as loop soft, $M_{\rm loop,soft}^{0\nu}$, so that
$M_{\rm usoft}^{0\nu}+$ $M_{\rm loop}^{0\nu}=\left(M_{\rm usoft}^{0\nu}+M_{\rm loop,usoft}^{0\nu}\right)$
$+\left(M^{0\nu}_{\rm loop,VV}+M^{0\nu}_{\rm loop,AA}+M^{0\nu}_{\rm loop,CT}\right)$
$=M_{\rm tot,usoft}^{0\nu}+M_{\rm loop,soft}^{0\nu}\,.$ (13)
Table 1: Central values and ranges of all NMEs up to N2LO: LO long range (second and sixth columns), LO short range (third and seventh), N2LO total ultrasoft (fourth and eighth) and N2LO soft loops (fifth and ninth). The NME ranges cover two interactions for the NSM or a range of $g_{pp}$ values for the pnQRPA, as well as various couplings $g_{\nu}^{\rm NN}$, Gaussian regulators $\Lambda=349\cdots 550$ MeV, CD-Bonn and Argonne SRCs, the renormalization scale range $\mu=500\cdots 1500$ MeV, and the ultrasoft scale $\mu_{\rm us}=m_{\pi}/2\cdots 2m_{\pi}$ (see text for details).
@@
L S tot usoft loop soft L S tot usoft loop soft
@@ 4.1 Ultrasoft NMEs and closure approximation
Table 1 presents the results of all NMEs up to N2LO calculated with the NSM and pnQRPA. The central values and ranges cover the interactions, couplings, regulators and SRCs discussed in Secs. 2 and  3. Focusing on the central values, the results in Table 1 indicate that for the NSM the total ultrasoft NME is about $5\%$-$10\%$ and has opposite sign compared to the total LO NME—summing the long- and short-range parts. The only exception is 48Ca, where the ultrasoft NME is suppressed. In contrast, for the pnQRPA the total ultrasoft NME has the same sign and $5\%$-$10\%$ of the value of the LO NME. Nonetheless, in 100Mo it reaches $40\%$ and it is very small in 116Cd and 136Xe. The uncertainty in $M_{\rm tot,usoft}^{0\nu}$ is significant, reaching $50\%$ for several nuclei in the NSM and pnQRPA calculations. This error is a measure of the inconsistency for not using a $\chi$EFT Hamiltonian, as it stems mainly from the $\mu_{\rm us}$ dependence of the NSM and pnQRPA NMEs. For the NSM the error in $M^{0\nu}_{\rm usoft}$ dominates, whereas for the pnQRPA the uncertainties for $M^{0\nu}_{\rm usoft}$ and $M^{0\nu}_{\rm loop,usoft}$ are comparable.
@@
Figure 1: Running sum of the total ultrasoft NME of 76Ge, $M_{\rm tot,usoft}^{0\nu}$, as a function of the excitation energy of the intermediate $1^{+}$ states, $E$, for the NSM and pnQRPA. The NSM results include the JUN45 (orange) and GCN2850 (red) interactions, while the pnQRPA results comprise $g_{\rm pp}$ values adjusted to $2\nu\beta\beta$-decay with $g_{\rm A}^{\rm eff}=1.0$ (blue) and $g_{\rm A}=1.27$ (cyan). The ranges cover the variation of the scale $\mu_{\rm us}$ from $m_{\pi}/2$ (dotted) to $2m_{\pi}$ (dashed).
In order to investigate the opposite sign of the ultrasoft NME in NSM and pnQRPA calculations, Fig. 1 compares the running sums of the $M_{\rm tot,usoft}^{0\nu}$ in 76Ge as a function of the excitation energy of the intermediate states, $E$, for both many-body methods. The results cover the NSM interactions and pnQRPA $g_{\rm pp}$ range presented in Sec. 3. Figure 1 shows that up to $E\approx 8$ MeV the two running sums are rather similar. However, they differ markedly around $8\text{ MeV}\lesssim E\lesssim 13\text{ MeV}$, where large cancellations and a resulting sign change appear in the pnQRPA but are not present in the NSM results. In fact, this behaviour is common to all the studied nuclei with the pnQRPA, while in the NSM we only observe $M_{\rm tot,usoft}^{0\nu}$ cancellations, although milder, in 48Ca. Similar differences between running sums in the NSM and pnQRPA have been noticed in the context of $2\nu\beta\beta$ decay [64]—which involves a similar sum over $1^{+}$ intermediate states. While the NSM may not capture these high-energy contributions because of missing spin-orbit partners in the valence spaces used, the pnQRPA may overestimate them. Measurements of charge-exchange reactions up to the GT resonance energy in both $\beta^{-}$ and $\beta^{+}$ directions, currently limited to lower energies [65, 66], could solve this discrepancy.
@@
In $\chi$EFT [36], $M^{0\nu}_{\rm usoft}$ can be seen as a correction beyond closure [67, 68, 69, 70, 71, 72, 73], since it is the lowest-order NME that explicitly depends on the structure of the intermediate states. We compare the difference of $M^{0\nu}_{\rm L}$ calculated with and without closure approximation, $\Delta_{\rm cl}\equiv M^{0\nu}_{\rm non-cl}-M^{0\nu}_{\rm L}$, with $M^{0\nu}_{\rm tot,usoft}$, which takes into account the uncertainty due to $\mu_{\text{us}}$ in $M^{0\nu}_{\rm usoft}$ and also in the related term $M^{0\nu}_{\rm loop,usoft}$. This uncertainty, reflected in the error bars, would vanish in a fully consistent calculation. We compute $\Delta_{\rm cl}$ in the pnQRPA and use the results of [71, 72, 73] for the NSM—for 130Te we average between the closure energies for 136Xe [72] and 124Sn [73]. Figure 2 shows the comparison. Considering the uncertainties, $M^{0\nu}_{\rm tot,usoft}$ is consistent with beyond-closure effects in all nuclei for the NSM. In contrast, for the pnQRPA, $M^{0\nu}_{\rm tot,usoft}$ amounts to roughly twice $\Delta_{\rm cl}$ in all nuclei, except for 136Xe where the two quantities are small. Nonetheless, the sign of both quantities is consistently the same.
@@
Figure 2: Difference between non-closure and LO long-range NMEs ($\Delta_{\text{cl}}$, pink and gray bars) compared with the total ultrasoft NME ($M_{\rm tot,usoft}^{0\nu}$, red and dark blue bars) obtained by the NSM and pnQRPA for all decays studied in this work. The NSM results use closure energies from [71, 72, 73] (solid bars) except for 130Te where we average the closure energies of 124Sn and 136Xe (open bar). Results for Argonne SRCs.
@@ 4.2 N2LO loop NMEs
Table 1 presents the calculated central values and ranges of the N2LO loop NMEs. While the central values are typically small, less than $5\%$ of the total LO NME for both many-body methods, the uncertainties are significant. When included, $M^{0\nu}_{\rm loop,soft}$ can reach up to $10\%$ of the total LO NME for the pnQRPA—$15\%$ for 100Mo—, and up to $5\%$ for the NSM, which are roughly the relative effects of the ultrasoft NMEs.
Figure 3: Radial distributions of the N2LO loop NME for 76Ge, $C^{0\nu}_{\rm loop,soft}$, obtained with the pnQRPA (left) and NSM (right). The results cover Gaussian regulators $\Lambda=349$ MeV (purple) and $\Lambda=550$ MeV (green), renormalization scales $\mu=500$ MeV (dark colors) and $\mu=1500$ MeV (light), ultrasoft scale $\mu_{\text{us}}=m_{\pi}$, and Argonne (dashed lines) and CD-Bonn (solid) SRCs.
To illuminate the origin of the large uncertainties for $M^{0\nu}_{\rm loop,soft}$, Fig. 3 shows, for the NSM and pnQRPA, their radial distribution for 76Ge, which satisfies $M^{0\nu}_{\rm loop,soft}=\int_{0}^{\infty}C^{0\nu}_{\rm loop,soft}(r){\rm d}r$. Different curves show the results for different regulators $\Lambda$, scales $\mu$, and SRCs. Qualitatively, the radial distributions look very similar in the pnQRPA (left) and NSM (right), but the overall scales are quite different. This is consistent with the comparison of the LO NMEs [39]. Figure 3 also indicates that the main sensitivity of the radial distribution of $M^{0\nu}_{\rm loop,soft}$ is to the regulator $\Lambda$ (difference between the blue and red curves: choosing a dipole regulator like in [43] gives results in between these two results). This sensitivity stems from the short-range—high-momentum—nature of the loop NME, also highlighted in Fig. 3, which concentrates at short-distances almost like the LO short-range NME [39]. This makes the N2LO loop NME very sensitive to SRCs as well: choosing either CD-Bonn or Argonne parameterizations can even change the sign of $M^{0\nu}_{\rm loop,soft}$. In contrast, the dependence on the scale $\mu$ is rather mild.
@@
124Sn $4.88$ $2.18$ $0.58$ $-0.49$ $0.19$ $0.85$ $3.17$ $0.90$ $-0.28$ $-0.11$ $0.05$ $0.24$
130Te $3.36$ $1.52$ $0.44$ $-0.39$ $0.15$ $0.67$ $3.24$ $0.94$ $-0.29$ $-0.12$ $0.06$ $0.19$
136Xe $2.71$ $0.99$ $0.11$ $-0.23$ $0.09$ $0.40$ $2.60$ $0.75$ $-0.23$ $-0.09$ $0.04$ $0.15$
Table 2 gives a decomposition of the VV, AA and CT components of the N2LO loops NME for a particular combination of nuclear interaction, couplings, regulators and SRCs—common for the NSM and pnQRPA results. The LO and N2LO ultrasoft NMEs obtained with the same combination of parameters are also listed for comparison. The results for the components of $M^{0\nu}_{\rm loop,soft}$ indicate that, both for the NSM and pnQRPA, the different parts cancel, with the AA and CT terms sharing the same sign as the LO NMEs, and VV carrying systematically the opposite sign. Individually, each of the three N2LO loops contributions of soft neutrinos can be of similar size, or even larger than the N2LO ultrasoft NME; however, due to the cancellation, the total $M^{0\nu}_{\rm loop,soft}$ tends to be smaller than $M^{0\nu}_{\rm tot,usoft}$, see Table 1.